The marginal cost of producing the Kindle is estimated at $126 per unit. Apply the MR = MC rule to find the output and price that maximize Kindle profits.



Considering that each Kindle sold generates $100 in e-book profits, determine Amazon’s optimal quantity and price with respect to the total profit generated by Kindle and e-book sales. What is the implication for Amazon’s pricing strategy?



1
Expert's answer
2022-03-08T12:40:40-0500

b)

"R=PQ\space \space AP+bQ^ 2"


"MR=a+2bQ"

"=294-2(0.000035)Q"


"MR=126"


"MC=126"


"Optimal \space Qty= 2,400,000.00"


"Price = 210.00"


c)

The MR of Kindle is effective "\$ \space 100 \space more"


"MR=394-2(0.000035) Q"


"MR=126"


"MC = 126"


"Optimal\space Qty=3,828,571.43"


"Price=160.00"


Amazon's optimal quality is higher.

Amazon can further decrease its price from the current 189 to 160